OpenStudy (anonymous):

Given that f(x) = 2x + 1 and g(x) = −5x + 2, solve for f(g(x)) when x = 3. −91 −33 −25 35 Fan & Medal for answer! :)

OpenStudy (anonymous):

Simply plug in 3 into g(x), and whatever you get for that answer, plug it in to f(x) and solve.

OpenStudy (anonymous):

I did that and got 4 different answers that weren't given.

OpenStudy (anonymous):

I got one of the answers. \[g(x)=-5(3)+2=-13 \] \[f(x)=2(-13)+1=-25\]

OpenStudy (anonymous):

But I thought they were set up differently?

OpenStudy (anonymous):

What do you mean? Since it's asking for f(g(x)), all you do is plug in 3 for g(x) and then you take the answer from that equation, and plug it into x in f(x).